OpenStudy (anonymous):

Solve the following system of equations 2x+3y-z=1 3x+y+2z=12 x+2y-3z=-5

OpenStudy (anonymous):

-6x-2y-4z=24 x+2y-32=-5 BECOMES -5x-7z=-29 2x+3y-z=1 -9x-3y-6z=-36 BECOMES 7x-7z=-36 bring back above equation *-1 5x+7z=29 BECOMES 12x=-7 x=-7/12 Plug this x value into 7x-7z=-36 to find z, then plug those values for x and z into one of the original equations to solve for the final variable

OpenStudy (anonymous):

so (3,1,2)?